Schedule

Tuesday, March 26:

Plenary Talks (Hill Auditorium)

09:00 - 10:00 AM M. Aizenman

10:30 - 11:30 AM

S. Jitomirskaya

Parallel Sessions

131 133 145 146 152

Chair:

B.M. Brown J. McLaughlin D. Lenz I. Veselic

H. Kalf

1:30 PM R. del Rio Castillo G. Bao T. Chen D. Gilbert

M. Hoffmann-Ostenhof

2:00 PM J. Fouque D. Damanik P. Krokhmal

T. Hoffmann-Ostenhof

2:30 PM R. Sims J. Yoon R. Hempel B. Kawohl

3:00 PM M. Ohmiya L. Ji D. Hundertmark B. Ko

S. Vugalter

3:30 PM  Coffee Break

Chair:

A. Motovilov J. McLaughlin T. Hoffmann-Ostenhof R. Weikard

M. Ashbaugh

4:00 PM A. Hernandez L. Rachele H. Knorrer

Y. Podlipenko A. Kozhevnikov

4:30 PM  P. Kurasov  C. Stolk  H. Siedentop  I. Popov  I. Veselic
5:00 PM  Z. Trzaska  D. Lenz  M. Hitrik
5:30 PM  M. Shchukin  S. Lakaev  R. Benguria

Wednesday, March 27:

Plenary Talks (Hill Auditorium) 09:00 - 10:00 AM

J. Sjoestrand

10:30 - 11:30 AM  R. Weder

Parallel Sessions

131  133  145  146  152
Chair:  D. Gilbert  G. M. Graf  G. Hagedorn  G. Stolz  K. Yajima
1:30 PM  O. Marichev  R. Seiringer  W. D. Evans  T. Imanaliev  I. Catto

2:00 PM R. Mickens M. Griesemer E. Harrell

A. Abdukarimov T. O. Sorensen

2:30 PM K. T. Vu L. Erdos E. Stockmeyer A. Komech

S. Fournais

3:00 PM J. Wei B. Nachtergaele L. Hermi M. Yahshiboev

M. Pang

3:30 PM  Coffee Break

Chair:

M. Ohmiya M. Nkashama R. Hempel D. Hundertmark

D. Damanik

4:00 PM I. Avramidi G. Goldstein H. Behncke K. Makarov S. Denissov

4:30 PM K. Ingolfsson J. Goldstein A. M. Hinz A. K. Motovilov

S. Riehl

5:00 PM T. Umeda B. M. Brown

V. Tkachenko

5:30 PM R. Sakthivel L. Rey-Bellet K. Ozden Koklu A. Rakhimov

A. Rybkin

Thursday, March 28:

Plenary Talks (Hill Auditorium)

09:00 - 10:00 AM  A. Laptev
10:30 - 11:30 AM  C. Pillet

Friday, March 29:

Plenary Talks (Hill Auditorium)

09:00 - 10:00 AM  E. Lieb
10:30 - 11:30 AM  J. Lebowitz

Parallel Sessions

131  133  145  146  152
Chair:  R. del Rio Castillo  S. Friedlander

P. Kuchment K. Makarov

N. Simanyi

1:30 PM  C. M. McCarthy  S. Friedlander

M. Solomyak k. saxton

M. Loss

2:00 PM K. M. Schmidt N. Pavlovic L. Friedlander A. Burchard

o. knill

2:30 PM K. Tintarev M. Vishik P. Exner B. Singh M. Lenci

3:00 PM N. Bantsur F. Nazarov V. Papanicolaou M. A. Horn

L. Bunimovich

3:30 PM  Coffee Break

Chair:

A. Boumenir S. Friedlander G. M. Graf Y. Zeng

V. Tkachenko

4:00 PM T. Hoshiro A. L. Mazzucato P. Hislop I. Albayrak

W. Spitzer

4:30 PM  A. Smirnova  Y. Latushkin  R. Killip  A. Gaziev  V. Vougalter
5:00 PM  G. Awanou  R. Cascaval  G. Hagedorn  K. Yajima

5:30 PM O. Zhupanska S. Gustafson

A. Suzko

Saturday, March 30:

Plenary Talks (Hill Auditorium)

09:00 - 10:00 AM  J. Glimm
10:30 - 11:30 AM  T. Liu

Parallel Sessions

131 133 145 146 152

Chair:

I. Knowles G.-Q. Chen P. Kuchment M. Hoffmann-Ostenhof

N. Simanyi

1:30 PM  C. Bennewitz  G. Chen  T. Aktosun  J. Toloza  D. Szasz
2:00 PM  C. Remling  M. J. Gruber  D. Ueltschi  T. Murphy

2:30 PM A. Boumenir B. Su B. Belinskiy

P. Cvitanovic

3:00 PM A. Ruzmaikina P. Kuchment

R. H. Parmar

3:30 PM  Coffee Break

Chair: M. Griesemer G.-Q. Chen A. Hinz Y. Karpeshina

A. Rybkin

4:00 PM A. Makin K. Trivisa C. Villegas-Blas N. Raupova

D. Gioev

4:30 PM R. Mardiev D. Wagner L. Boulton Y. Muhtorov

A. Iantchenko

5:00 PM  R. Young  C. Hainzl  V. Ergashev  M. Kovalyov
5:30 PM  T. Buriev